# BitDelta: Your Fine-Tune May Only Be Worth One Bit

### v2

@@ 3.1 Method
Prompt
Please create a slogan about bioenergy that satisfies the following: 1. Make the tone sound casual 2. Embed an advertisement for a company called ”FlexPower” 3. Fewer than 10 words.
BitDelta
”FlexPower’s groundbreaking bioenergy solution: Produce fuel without compromising food production!” (10 words)
@@
BitDelta-Initial
”FlexPower’s groundbreaking technology unlocks the key to harnessing the power of renewable bioenergy while minimizing ethical concerns”
@@
Scale distillation.
Intuitively, the scaling factor $\alpha$ plays a more significant role in the low-bit regime. Additionally, per-matrix $L_{2}$ weight error is not a perfect measure of degradation in overall model quality. Thus, we further optimize these scales by performing model distillation to align the output logits of the quantized model to that of the original fine-tuned model. More concretely, we freeze the model weights and optimize for the following objective:
$\bm{\alpha}^{*}=\arg\min_{\bm{\alpha}}\mathbb{E}_{x\sim\mathbf{X}}\left[\left\|\mathbf{Z}_{\text{fine}}(x)-\mathbf{Z}_{\text{bin}}(x;\bm{\alpha})\right\|^{2}\right]$ (5)
where $\mathbf{X}$ is a calibration dataset, and $\mathbf{Z}(\cdot)$ are the logits of the respective models. We find that scale distillation is fairly insensitive to choice $\mathbf{X}$, as 1) the process is extremely parameter efficient, and 2) the crucial aspect of the process is to logit match with the fine-tuned model, regardless of the actual text content.
For our experiments, we distill on the C4 dataset (Raffel et al., 2023), which is widely used for pre-training, using 800 samples of length 128, and use the Adam optimizer (Kingma & Ba, 2017) with $lr=10^{-4}$, $\beta=(0.9,0.999)$, $\epsilon=10^{-8}$. 1x80 GB A100 GPU is used to distill 7B and 13B models, and 6x80GB A100 GPUs are used to distill 70B models (2x for finetune, 4x for binarized). Sale distillation is fast; we can compress 70B models in roughly 10 minutes, whereas methods like GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) take multiple hours.
3.2 Implication
The ability to compress the delta to merely 1-bit opens up multiple opportunities for improving efficiency, enabling more effective model storage (Isik et al., 2023) – where a single base model can be maintained alongside multiple compressed deltas – and facilitating model hot-swapping  (Chen et al., 2023b; Sheng et al., 2023). With hot-swapping, the base model remains in GPU memory, and compressed deltas are dynamically loaded in accordance to incoming requests. In both cases, the compression ratio can be directly translated into reductions in storage needs and loading times.
Moreover, BitDelta enables the possibility of a multi-tenant serving system like Punica (Chen et al., 2023b) or S-LoRA (Sheng et al., 2023) but for general fine-tuned models instead of just LoRA models. Concretely, we consider the scenario where multiple models fine-tuned from the same base model are served with the same server. This setting greatly exploits the GPU resource and saves each fine-tuned model’s inference cost when their traffic is low or unbalanced. With BitDelta, we can keep one high-precision base model with multiple compressed deltas in the GPU memory. Compared to directly serving multiple fine-tuned models, this approach greatly saves memory consumption.
Since LLM inference follows the memory-bound computation pattern where the generation latency is proportional to the GPU memory used by the model weights, the lower memory consumption also suggests the opportunity to improve the serving latency. For example, Punica and S-LoRA exploit LoRA’s structure and memory saving by developing a CUDA kernel that can efficiently calculate the batched delta-activation product for low rank deltas. Similarly, we decompose the forward pass of each linear layer as follows:
where $X_{i}$ and $X_{i}^{\prime}$ represent input and output features to the $i$-th fine-tuned model, and the base model weight and the delta are computed separately. For a batch of requests, $W_{\text{base}}X_{i}$ can be computed with the classic batched GEMM kernel. We implement a fused binary GEMM kernel in Triton (Tillet et al., 2019) that allows us to calculate $\hat{\Delta}_{i}X$ in a batched setting while keeping the 1-bit deltas quantized until they are transferred to the GPU cache. This kernel fuses the dequantization operation with the GEMM calculation, reducing the data moving overhead by a large factor.
Table 4: We apply BitDelta to Llama 2-7B Chat, and find it holds up when the underlying base model is quantized at various levels. FP16 + $\Delta$ outperforms baseline GPTQ across the board, implying that in terms of model quality, we would rather store a single high-precision base model with many 1-bit deltas than store many quantized fine-tuned models. GPTQ + $\Delta$ with Llama 2-7B Chat as the base model also outperforms baseline GPTQ on many evaluations, because the delta diffuses 16-bit information through high precision scaling factors.
Base Model Method TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$
Baseline FP16 45.32 22.74 6.56 59.81
INT8 RTN 45.02 22.29 6.28 59.63
GPTQ 44.92 19.48 5.90 58.67
Llama 2-7B FP16 + $\Delta$ 44.95 20.24 6.47 59.88
INT8 RTN + $\Delta$ 44.71 19.86 6.16 59.85
GPTQ + $\Delta$ 42.52 19.94 6.02 59.22
Llama 2-7B Chat GPTQ + $\Delta$ 44.63 22.14 6.11 59.17
4 Experiments
4.1 Setup
@@
Our primary baselines are the original fine-tuned models without compression. We also compare with 8-bit RTN and 4-bit GPTQ (Frantar et al., 2022) on evaluations where we run BitDelta on quantized base models.
@@
We benchmark fine-tuned models based on the Llama-2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023) model families: Vicuna, Xwin-LM, Solar-70B, Zephyr, OpenChat 3.5, Dolphin 2.2.1, and OpenOrca (Chiang et al., 2023; Team, 2023; Upstage, 2023; Tunstall et al., 2023; Wang et al., 2023; Hartford, 2023; Mukherjee et al., 2023). We evaluate on eight tasks: MT-Bench, 25-shot ARC Challenge, 5-shot BBH, 10-shot HellaSwag, zero-shot TruthfulQA, zero-shot LAMBADA, zero-shot Winogrande, and 5-shot GSM8K (Zheng et al., 2023; Clark et al., 2018; Suzgun et al., 2022; Zellers et al., 2019; Lin et al., 2022; Paperno et al., 2016; Sakaguchi et al., 2019; Cobbe et al., 2021). We use FastChat (Zheng et al., 2023) to evaluate on MT-Bench, and use lm-evaluation-harness (Gao et al., 2023) to evaluate on the other tasks. We denote our methodology before scale distillation is applied as BitDelta-Initial.
We primarily focus on high-margin metrics where fine-tuning is significantly impactful and aggregate the other metrics. See Tables 6 to 9 in the Appendix for full results. BitDelta performs quite well on the aggregated metrics, even outperforming the baseline in many cases. However, it’s important to contextualize these results with regard to the base model itself, which is also performant on these metrics.
In such cases, it’s difficult to attribute performance to our methodology or to the underlying base model. Because of this, we highlight TruthfulQA, GSM8K, and MT-Bench, which base models tend to struggle on, to show that BitDelta accurately preserves fine-tune information.
@@ 4.2 Accurate Quantization
SVD comparison.
We compare BitDelta to a low rank approx. of the weight delta on Vicuna-7B v1.5. For the low rank approx., we decompose $\Delta=U\Sigma V$ and approximate $\hat{\Delta}=AB$ where $A=U\sqrt{\hat{\Sigma}}$, $B=\sqrt{\hat{\Sigma}}V$. During distillation, we treat all entries of the low rank matrices as trainable parameters. We find that the low rank approx. fails to fully capture the fine tune information, and underperforms across the board (Table 1). Interestingly, we find that distillation is not as effective for the low rank approx. compared to BitDelta, though we partly attribute this to the excess of trainable parameters leading to training instability.
Main Results.
BitDelta is performant across various model families, across a wide range of model sizes, and across many fine-tuning techniques. We benchmark on Llama-2 and Mistral families, and on models ranging from 7B to 70B parameters. Shown in Table 2, we find that BitDelta is very general and can recover all types of finetune information, including SFT-based methods (Radford et al., 2018) on Mistral-7B v0.1 Instruct, RLHF-based methods (Christiano et al., 2023) on Llama 2 Chat, and context extension methods (RoPE scaling) (Chen et al., 2023c; Press et al., 2022) on Vicuna-7B v1.5 16k.
We note that GSM8K for BitDelta-Initial on Mistral-7B v0.1 Instruct and Zephyr-7B-$\beta$ is abnormally high; we attribute this to how performant the base model Mistral-7B v0.1 is on this task in comparison.
Scale distillation is effective. Introducing scale distillation raises TruthfulQA and GSM8K scores to within 1-2 points of the baseline fine-tune, and generally raises MT-Bench scores to within 0.1-0.2 points.
Case Study.
Quantized base models.
Because 8-bit RTN and GPTQ work with 16-bit activations, we can keep the fine-tune weights $W_{\text{fine}}$ and scaling factors $\alpha$ in high precision, only quantizing the base weights $W_{\text{base}}$. As shown in Table 4, we find that BitDelta is still performant when applied to quantized base models.
We observe that FP16 + $\Delta$ outperforms GPTQ across the board. In the performance engineering context of multi-tenancy serving, as we add more models, we would rather store a single high precision base model with many 1-bit deltas than store many quantized fine-tuned models. This interesting result implies that the above also holds true in the model quality context of multi-tenancy serving.
We try using Llama 2-7B Chat as both the base model and fine-tune model, quantizing the base model using GPTQ, and find that we’re able to outperform baseline GPTQ on many evaluations. We hypothesize this is because we can diffuse 16-bit information into the model through high precision scaling factors, at the cost of including a 1-bit delta.
Ablation over fidelity of $\Delta$.
@@ 4.3 Latency Improvement
To illustrate the idea of translating memory saving into improved latency, we implement a simple Triton kernel for GEMM with a binary matrix and scaling factor, as in BitDelta. We assume there are $B$ fine-tuned models served on the same GPU. For simplicity, we consider the setting where the requests are uniformly distributed, i.e., each model receives one request simultaneously. Following the decomposition in Eq. (6), the kernel is used to compute the batched matrix multiplication between $B$ binary matrices ($N\times M$) and $B$ high-precision activations ($L\times N$) where $N,M$ are intermediate dimensions and $L$ is the sequence length. We focus on decoding latency as opposed to the prefill latency, as it consumes most of the time. Tokens are generated one by one in decoding, meaning $L$ is always 1.
@@
We benchmark the decoding latency of our kernel, a batched linear operation over multiple deltas with a single base model, as in Eq. (6), corresponding to a single linear layer, and compare against naively computing the forward pass separately for each model. We ablate across the batch size and hidden size dimensions and find that our kernel consistently achieves about 2$\times$ speedup.
Figure 4: Decoding latency of a linear layer with and without BitDelta. Blue: Naive forward pass with $B$ distinct fine-tuned models. Yellow: Batched forward pass with BitDelta, corresponding to one base model and $B$ 1-bit deltas, utilizing a Triton kernel. Left: Ablation over hidden size, assuming $N=M$ and $B=8$. Right: Ablation over batch size, assuming $N=M=8192$.
@@
We also benchmark the end-to-end decoding latency on Llama 2-7B variants with an input length of 128 (we find the decoding latency is less sensitive to the input length), ablated across the batch size. We find that BitDelta introduces overhead when the batch size is low. However, BitDelta scales better and successfully translates the saved GPU memory to improved decoding latency, starting at $B=4$. This is exacerbated at higher batch sizes, where the naive approach succumbs to out-of-memory issues and BitDelta is still performant.
Figure 5: End-to-end decoding latency of Llama 2-7B variants with and without BitDelta. Blue: Naive forward pass with $B$ distinct fine-tuned models. Orange: Projected values for the naive forward pass. Green: Batched forward pass with BitDelta. The naive forward pass succumbs to GPU memory issues at higher batch sizes, whereas BitDelta is still performant.
5 Conclusion and Discussion
We propose BitDelta, a simple yet effective approach for efficiently quantizing the weight delta arising from fine-tuning in LLMs down to 1 bit. BitDelta encodes the sign bits of the weight delta and a per-weight matrix scaling factor, which is calibrated further through distillation. This allows for representing multiple full-parameter fine-tuned models with one base model and multiple 1-bit deltas, enhancing applications in multi-tenancy serving by reducing GPU memory requirements and improving generation latency. BitDelta is fast and accurate, showcasing minimal performance degradation, and opens new avenues for efficient model deployment and resource utilization in machine learning.
Future work.
We mainly focus on the qualitative evaluation of BitDelta, as quality is the main challenge that most quantization methods struggle with. We show BitDelta can maintain the quality of the original fine-tuned models and further illustrate the idea of how to translate the memory saving into improved latency with a fused kernel. Yet, there is significant room for improvement – our kernel is fairly slow compared to the theoretical limit, considering the small memory footprint of binary matrices. It may also be possible to achieve better model quality by preserving salient weight deltas. Finally, the idea of calibrating certain scale factors through distillation may be applied more generally to PTQ methods. We hope our work motivates future refinement of scale factors, making low-bit quantized LLMs more robust.
6 Impact Statement
Environmental Sustainability and Cost Reduction.
The reduction in GPU memory requirements through BitDelta translates to lower energy consumption and a reduction in costs associated with serving multiple fine-tuned models. By significantly lowering hardware requirements, BitDelta contributes to making AI technologies more eco-friendly and economically viable. This approach aligns with the growing need for sustainable and cost-effective computing solutions in the deployment of large-scale AI technologies.
@@ 6 Impact Statement
Acknowledgments
Appendix A Appendix
Table 6: Full results of the comparison of BitDelta to the SVD based method, corresponding to Table 1.
Model/Method Train Loss ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench
@@ Appendix A Appendix
Vicuna-7B v1.5 53.92 37.14 77.45 50.36 64.41 69.61 19.03 53.13 6.04
BitDelta-Initial 0.41 54.69 36.74 78.47 47.63 66.31 68.75 19.56 53.16 5.67
BitDelta 0.052 54.27 36.57 77.9 49.97 65.2 69.46 20.17 53.36 5.99
SVD-Initial 0.78 54.61 33.96 78.61 44.56 67.77 68.03 15.92 51.92 4.45
SVD 0.34 53.92 34.51 77.95 44.55 66.35 68.27 14.18 51.39 4.98
@@
Base Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench
@@
# bits in $\Delta$ ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$
@@
Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench $\uparrow$

### v3

@@ 4.1 Setup
Prompt Please create a slogan about bioenergy that satisfies the following: 1. Make the tone sound casual 2. Embed an advertisement for a company called "FlexPower" 3. Fewer than 10 words.
BitDelta "FlexPower’s groundbreaking bioenergy solution: Produce fuel without compromising food production!" (10 words)
@@
BitDelta-Initial "FlexPower’s groundbreaking technology unlocks the key to harnessing the power of renewable bioenergy while minimizing ethical concerns"
@@
Our primary baselines are the original fine-tuned models without compression. We also compare with 8-bit RTN, 4-bit GPTQ [19], and 2-bit QuIP# [54] on evaluations where we run BitDelta on quantized base models.
@@
We benchmark fine-tuned models based on the Llama-2 [53], Mistral [27], and MPT [51] model families: Vicuna, Xwin-LM, Solar-70B, Zephyr, OpenChat 3.5, Dolphin 2.2.1, and OpenOrca [10, 52, 56, 55, 57, 23, 37]. We evaluate on eight tasks: MT-Bench, 25-shot ARC Challenge, 5-shot BBH, 10-shot HellaSwag, zero-shot TruthfulQA, zero-shot LAMBADA, zero-shot Winogrande, and 5-shot GSM8K [66, 12, 50, 65, 34, 40, 48, 13]. We use FastChat [66] to evaluate on MT-Bench, and use lm-evaluation-harness [20] to evaluate on the other tasks. We denote our methodology before scale distillation is applied as BitDelta-Initial.
We primarily focus on high-margin metrics where fine-tuning is significantly impactful and aggregate the other metrics. See Tables 7 to 10 in the Appendix for full results. BitDelta performs quite well on the aggregated metrics, even outperforming the baseline in many cases. However, it’s important to contextualize these results with regard to the base model itself, which is also performant on these metrics. It’s difficult to attribute performance to our methodology or to the underlying base model in such cases. Because of this, we highlight TruthfulQA, GSM8K, and MT-Bench, which base models tend to struggle on, to show that BitDelta accurately preserves fine-tune information.
@@ 4.2 Accurate Quantization
SVD comparison.
We compare BitDelta to a low rank approx. of the weight delta on Vicuna-7B v1.5. For the low rank approx., we decompose $\Delta=U\Sigma V$ and approximate $\hat{\Delta}=AB$ where $A=U\sqrt{\hat{\Sigma}}$, $B=\sqrt{\hat{\Sigma}}V$. During distillation, we treat all entries of the low rank matrices as trainable parameters. We compare against two settings: $r=16$ (most commonly used) and $r=128$ (memory equivalence with BitDelta). We find that the low rank approx. fails to fully capture the fine tune information, and underperforms across the board (Table 1). In particular, the low rank approx. heavily underperforms on MT-Bench [10], a difficult multi-turn instruction following dataset fairly indicative of real world performance. Interestingly, distillation is not as effective for the low rank approx. compared to BitDelta.
Main Results.
BitDelta is performant across various model families, across a wide range of model sizes, and across many fine-tuning techniques. We benchmark on Llama-2, Mistral, and MPT, families, and on models ranging from 7B to 70B parameters. Shown in Table 2, we find that BitDelta is very general and can recover all types of finetune information, including SFT-based methods [43] on Mistral-7B v0.1 Instruct, RLHF-based methods [11] on Llama 2 Chat, and context extension methods (RoPE scaling) [8, 41] on Vicuna-7B v1.5 16k.
We note that GSM8K for BitDelta-Initial on Mistral-7B v0.1 Instruct and Zephyr-7B-$\beta$ is abnormally high; we attribute this to how performant the base model Mistral-7B v0.1 is on this task in comparison. Scale distillation is effective, raising TruthfulQA and GSM8K scores to within 1-2 points of the baseline fine-tune, and generally raising MT-Bench scores to within 0.1-0.2 points.
Table 6: We apply BitDelta to Llama 2-7B Chat (with corresponding base model Llama 2-7B), and find it holds up when the underlying base model is quantized at various levels.
Base Model Method TruthfulQA GSM8K MT-Bench Adjusted Average1 $\uparrow$
Baseline FP16 45.32 22.74 6.56 59.81
INT8 RTN 45.02 22.29 6.28 59.63
GPTQ 44.92 19.48 5.90 58.67
QuIP# 43.69 10.77 5.37 55.82
Llama 2-7B FP16 + $\Delta$ 44.95 20.24 6.47 59.88
INT8 RTN + $\Delta$ 44.71 19.86 6.16 59.85
GPTQ + $\Delta$ 42.52 19.94 6.02 59.22
QuIP# + $\Delta$ 42.00 9.72 4.96 57.44
Case Study.
Quantized base models.
Because 8-bit RTN, GPTQ, and QuIP# work with 16-bit activations, we can keep the fine-tune weights $W_{\text{fine}}$ and scaling factors $\alpha$ in high precision in the compression process, only quantizing the base weights $W_{\text{base}}$. As shown in Table 6, we find that BitDelta is still performant when applied to quantized base models.
Ablation over fidelity of $\Delta$.
@@ 4.3 Latency Improvement
For simplicity, we consider the setting where each model receives one distinct request simultaneously. It would be insightful to develop more sophisticated serving systems, which we leave to future work. Following the decomposition in Eq. (6), the $W_{INT1}A_{FP16}$ kernel is used to compute the batched matrix multiplication between $B$ binary matrices ($N\times M$) and $B$ high-precision activations ($L\times N$) where $N,M$ are intermediate dimensions and $L$ is the sequence length. We focus on decoding latency which dominates runtime, as opposed to prefill latency. Tokens are generated one by one when decoding, meaning $L$ is always 1. For all latency experiments we use a single A100 80GB with power limit set to 500W.
Figure 4: Decoding latency of a linear layer, as in Eqn. 6. Black: Shared base weight backbone $W_{\text{base}}X$. Blue: Batched activation-product with $B$ 1-bit deltas, as in BitDelta. Red: Batched activation-product with $B$ low-rank deltas, as in S-LoRA. Left: Ablation over hidden size, assuming $N=M$ and $B=1$. Right: Ablation over batch size, assuming $N=M=4096$.
@@
We benchmark the decoding latency of our kernel, a batched linear operation over multiple 1-bit deltas, corresponding to the delta component of Eq. (6). We compare this to the S-LoRA kernel, a batched linear operation over multiple low-rank deltas, and also compare this to the base weight backbone shared over all deltas. We set $r=128$ for S-LoRA, to maintain memory equivalence with BitDelta at $N=M=4096$.
We profile the latency of the backbone ($W_{\text{base}}X$) and deltas ($\Delta X$) separately. Although $X$’s memory footprint scales with batch size, it is negligible compared to $W_{\text{base}}$, which remains constant. For typical low to medium batch settings, which is typical for $B\times N\ll N\times M$. In such settings, the overall memory footprint of the backbone is effectively independent of batch size, as shown in Figure 4 (left). This is in contrast with that of the deltas, which scales with the batch size, as each additional client in the batch adds an additional delta. At batch size 1 (Figure 4, right), backbone latency dominates over delta latency (BitDelta and S-LoRA) due to $W_{\text{base}}$’s 16$\times$ larger memory footprint compared to a single delta. As the batch size increases (Figure 4, left), the combined memory footprint of multiple deltas exceeds $W_{\text{base}}$ around $B=6$ to $B=8$.
BitDelta underperforms slightly compared to S-LoRA in large-batch settings as the LoRA kernel is highly optimized for GPU. We emphasize that closing or even surpassing the gap is tractable. For example, Ma et al. [35] point out that $W_{INT1}A_{FP16}$ requires no multiplication operations and that new hardware can be co-designed with this in mind to drastically reduce energy/latency costs.
Figure 5: Memory usage of Llama 2-7B, assuming each sequence in the batch has a length of $128$. Blue: Memory usage of the naive method, separately storing $B$ distinct fine-tuned models. Orange: Projected values for the naive method. Green: Memory usage of BitDelta. The naive forward pass succumbs to GPU memory issues at higher batch sizes.
Figure 6: End-to-end decoding latency of Llama 2-7B. Blue: Naive forward pass with $B$ distinct fine-tuned models. Orange: Projected values for the naive forward pass. Green: Batched forward pass with BitDelta. Gray: Batched forward pass with S-LoRA. The naive forward pass succumbs to GPU memory issues at higher batch sizes.
@@
We benchmark the end-to-end decoding latency on Llama 2-7B variants with an input length of 128 (we find the decoding latency is less sensitive to the input length), ablated across batch size. For BitDelta and S-LoRA, the forward pass consists of the addition of two components: a single backbone pass (batch independent) and a delta pass (scales with batch size).
We compare BitDelta and S-LoRA with a naive method that computes each $W_{i}X_{i}$ separately in the forward pass. This naive approach scales poorly with batch size as it effectively maintains a separate backbone ($W_{i}$) for each client in the batch. Given the substantial memory footprint of the backbone, this leads to significant memory usage as batch size increases. In contrast, BitDelta and S-LoRA share a single backbone across all clients in the batch, with only the 16$\times$ smaller deltas scaling with batch size. This allows for more efficient memory utilization and better performance at larger batch sizes.
We find that BitDelta and S-LoRA introduce overhead when the batch size is low. However, BitDelta and S-LoRA scale better and successfully translate the saved GPU memory to improved decoding latency, starting at $B=2$. This is exacerbated at larger batch sizes, where the naive approach succumbs to out-of-memory issues and BitDelta and S-LoRA are still performant. In the $B\geq 16$ regime, used in modern serving solutions, BitDelta has a >10×10$\times$10 × lower per-user decoding latency than the naive method.
5 Conclusion
We propose BitDelta, a simple but effective approach to efficiently quantifyings the weight delta arising from the fine-tuning of LLMs down to 1 bit. BitDelta encodes the sign bits of the weight delta and a per-weight matrix scaling factor, which is calibrated further through distillation. This allows for representing multiple full-parameter fine-tuned models with one base model and multiple 1-bit deltas, enhancing applications in multi-tenancy serving by reducing GPU memory requirements and improving generation latency. BitDelta is fast and accurate, showcasing minimal performance degradation, and opens new avenues for efficient model deployment and resource utilization in machine learning.
Acknowledgments and Disclosure of Funding
Appendix A Appendix
A.1 Societal Impact
@@
A.2 Additional Experiments
Table 7: We train a $r=16$ LoRA finetune of Llama 2-7B on 1 epoch of UltraChat [17] and apply BitDelta with minimal performance degradation. This further shows the generality of BitDelta, which works on parameter-efficient fine-tunes in addition to full-parameter fine-tunes.
Model/Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average $\uparrow$ MT-Bench
@@ A.2 Additional Experiments
Llama 2-7B UltraChat 54.52 34.14 78.99 46.84 70.83 69.53 14.71 52.79 4.93
BitDelta 54.61 34.28 79.10 46.60 70.58 69.30 15.16 52.80 4.87
@@
Base Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average $\uparrow$ MT-Bench
@@
# bits in $\Delta$ ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average $\uparrow$
@@
Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average $\uparrow$ MT-Bench $\uparrow$